Skip to main content
A-LevelMathematicsProbability and statistics9709/52 · Feb/Mar 2025 · Q59 Marks

The mass of peaches sold per day in a supermarket is normally distributed with mean 65.8kg and standard deviation 9.6kg.

  1. (a)Find the probability that the mass of peaches sold on any given day is between 56kg and 75kg.[3]
  2. (b)The mass of cherries sold per day in a supermarket is normally distributed with mean 72.4kg and standard deviation σkg. It is known that on 10% of days less than 59.1 kg of cherries are sold. Find the value of σ.[3]
  3. (c)The supermarket is open 7 days a week. Find the probability that, in a randomly chosen week, the first day on which less than 59.1kg of cherries are sold is the fifth day of the week.[1]
  4. (d)Find the probability that, in a randomly chosen week, the first day on which less than 59.1kg of cherries are sold is before the fifth day of the week.[2]

Mark scheme

Marking points from the official mark scheme (Feb/Mar 2025). The mark or mark type is shown in brackets.

  1. [1] Use of standardisation formula once with 65.8, 9.6 and either 56 or 75. No continuity correction, not σ², not √σ.
  2. [1] Appropriate probability area Φ, from final process. Must be a probability. (expect > 0.5).
  3. [1] AWRT. If 1 or more M mark not awarded, SC1 for final answer 0.677 AWRT WWW.
  4. P(56 < X < 75) = P((56-65.8)/9.6 < Z < (75-65.8)/9.6) = P(-1.0208 < Z < 0.9583) = Φ(0.9583) - Φ(-1.0208) = Φ(0.9583) - (1 - Φ(1.0208)) = 0.8309 - (1 - 0.8463) = 0.8309 - 0.1537 = 0.6772. Rounded to 3 significant figures, 0.677.
  5. [1] 1.282 or -1.282 seen cao (critical value).
  6. [1] Standardisation formula with 59.1, 72.4, σ equating to a z-value (not 0.1, 0.9, 0.5398, 0.4602, 0.8159, 0.1841). Condone continuity correction of ±0.05, not σ² and not √σ. Condone ±13.3/σ = -1.282.
  7. [1] AWRT. Signs must be consistent throughout. If M1 not awarded, SC1 σ=10.4 WWW.
  8. P(X < 59.1) = 0.10. P(Z < (59.1 - 72.4)/σ) = 0.10. From tables, Φ(-1.282) = 0.10. So, (59.1 - 72.4)/σ = -1.282. -13.3/σ = -1.282. σ = 13.3 / 1.282 ≈ 10.374. Rounded to 3 significant figures, σ = 10.4.
  9. [1] [(0.9)^4 * (0.1) =]0.0656[1]
  10. [1] 1-(0.9)^d d = 4, 5.
  11. [1] 0.3439.
  12. P(X < 5) = 1 - P(X ≥ 5) = 1 - (0.9)^4 = 1 - 0.6561 = 0.3439. Rounded to 3 significant figures, 0.344.

Write your answer and get it marked

On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.

Sign up free →

About this question

Topic
Probability and statistics, A-Level Mathematics (syllabus 9709). Structured question, 9 marks.
Source
A-Level Mathematics, Feb/Mar 2025 Paper 5 Variant 2, question 5.

Questions and mark schemes are from past examination papers and remain the copyright of the exam board. How Oracle Prep builds its content

Related Mathematics questions

© 2026 Oracle Prep